OpenStudy (anonymous):

#HELP!! A company builds computers. It costs $6,700 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? C(x) = 550x −1200 C(x) = 550x + 1200 C(x) = 1200x − 550 C(x) = 1200x + 550

OpenStudy (valpey):

This is a fixed/variable cost problem. There is some fixed cost to running the place and some variable cost for each computer built so it should look like C(x) = x*VC + FC

OpenStudy (valpey):

$6,700 = 10*VC + FC $12,200 = 20*VC + FC

OpenStudy (anonymous):

UHH

OpenStudy (valpey):

Subtract the first equation from the second equation.